11 11 TEX L^{A}TEX

Linux VAR HOWTO

Mr. Poet poet@linuxports.com v2.84, 22 August 1999 This document contains a listing of service companies which do not manufacture hardware or create packaged software, but add value to existing products. It is maintained by Mr. Poet poet@linuxports.com.

Contents

1 About this document 4 2 Copyright Information 4 3 VAR ID: Australia, Cybersource 5 4 VAR ID: Australia, EMUSYS Unix Consulting 5 5 VAR ID: Australia, Sydnet Group Pty Ltd 6 6 VAR ID: Austria, QUANT-X Service & Consulting GmbH 7 7 VAR ID: Brazil, Millennium Consultoria e Informática Ltda. 7 8 VAR ID: Canada, Affinity Systems Inc. 8 9 VAR ID: Canada, Centurion Services 8 10 VAR ID: Canada, Mediabase 9 11 VAR ID: Canada, Net Direct Inc. 10 12 VAR ID: Canada, Softouch Computer Solutions 10 13 VAR ID: Canada, VL 11 14 VAR ID: Finland, SOT Finnish Software Engineering Ltd. 11 15 VAR ID: France, Paralline 12 16 VAR ID: Germany, IGEL GmbH 13 17 VAR ID: Germany, IGEL 13 18 VAR ID: Germany, dynamis EDV-Consulting 14 19 VAR ID: Hong Kong, Genesis Systems Int’l Ltd. 14 20 VAR ID: India, G.T.Enterprises 15 21 VAR ID: Indonesia, PT Cakram DataLingga Duaribu (CDL2000) 16 22 VAR ID: Italy, Conecta Telematica srl 16 23 VAR ID: Italy, Nauta Srl - Internet per l’impresa 17 24 VAR ID: Italy, Studio C 17 25 VAR ID: Japan, SoftWare Research Co. 18 26 VAR ID: Malaysia, Linux Resources Sdn. Bhd. 19 27 VAR ID: Malaysia, MAGNIFIX 19 28 VAR ID: Malaysia, Malaysian Linux Competency Center Sdn. Bhd. 20 29 VAR ID: Netherlands, CMG Trade, Transport & Industry 20 30 VAR ID: Netherlands, X OS Experts in Open Systems BV 21 31 VAR ID: Poland, Emit Sp z o.o. 22 32 VAR ID: Romania, EVERCOMM Serv SRL 22 33 VAR ID: Spain, Innovacion y Futuro s.l. 23 34 VAR ID: Spain, LANGRE INGENIEROS S.L. 23 35 VAR ID: Sweden, South Pole AB 24 36 VAR ID: Taiwan, Whinstone Info. Corp. 25 37 VAR ID: Turkey, Gelecek Bilisim ve Iletisim A.S 25 38 VAR ID: US, AZ, Sunset Systems 26 39 VAR ID: US, CA, ASL Workstations, Inc. 26 40 VAR ID: US, CA, Linux Laptops Ltd. 27 41 VAR ID: US, CA, PSSC Labs 28 42 VAR ID: US, CO, Apache Digital Corporation 28 43 VAR ID: US, CT, GNUTEC 29 44 VAR ID: US, FL, eBuilders Network 29 45 VAR ID: US, IL, Advanced Horizons, Inc 30 46 VAR ID: US, IL, QLITech Linux COmputers 31 47 VAR ID: US, IL, QLITech Linux Computers 31 48 VAR ID: US, IL, WellThot Inc. 32 49 VAR ID: US, KY, IndyBox Systems, Inc. 32 50 VAR ID: US, KY, Neville Technologies 33 51 VAR ID: US, MA, Network Concierge Inc. 34 52 VAR ID: US, MD, AmNet Computers 34 53 VAR ID: US, MN, Cranston Software 35 54 VAR ID: US, MN, Donnal Group Inc 35 55 VAR ID: US, MO, Atipa Linux Solutions 36 56 VAR ID: US, MO, Omnitec Corporatin 37 57 VAR ID: US, NC, ApeX Systems Integration Corp. 37 58 VAR ID: US, NC, ComputerNet Inc. 38 59 VAR ID: US, NC, Indelible Blue 38 60 VAR ID: US, NE, BrainStorm Technologies 39 61 VAR ID: US, NE, BrainStorm Technologies 40 62 VAR ID: US, NH, DCG Computers, Inc. 40 63 VAR ID: US, NON United States, Microway 41 64 VAR ID: US, NON United States, Telenet System Solutions, Inc. 41 65 VAR ID: US, NY, Logicept Corporation 42 66 VAR ID: US, OH, 800Linux 43 67 VAR ID: US, OH, LANshark Systems, Inc. 43 68 VAR ID: US, OH, The Computer Underground 44 69 VAR ID: US, OH, Unique Systems, Inc. 44 70 VAR ID: US, OR, 3D Technology, Inc. 45 71 VAR ID: US, PR, OG Consulting 45 72 VAR ID: US, TX, Computer Innovations International 46 73 VAR ID: US, TX, Heights Open Systems 47 74 VAR ID: US, TX, Parachute 47

75 VAR ID: US, VA, NetSource of Virginia, Inc. 48

76 VAR ID: US, WA, Cheek Consulting 48

77 VAR ID: US, WA, Lodgepole Technology 49

78 VAR ID: United Arab Emirates, JM Infotech Solutions STPI 50

79 VAR ID: United Kingdom, DGC-NMS 50

80 VAR ID: United Kingdom, Definite Linux Systems Ltd 51

81 VAR ID: United Kingdom, Digital Networks 51

82 VAR ID: United Kingdom, iCode Systems Ltd 52

1 About this document

This is the Linux VAR HOWTO. It contains a listing of service companies which do not manufacture hardware or create packaged software, but add value to existing products. This added value can take many forms, such as installation of Linux on computer hardware, configuration, or custom applications. However, at the moment the Linux VAR HOWTO only contains companies selling preconfigured hardware for the purpose of running Linux. Companies offering other services are invited to contact the maintainer poet@linuxports.com at poets@linuxports.com for a listing.

If you contact any companies listed in this document, please mention the Linux VAR HOWTO.

If you need to know more about the Linux Documentation Project or about Linux HOWTO’s, feel free to contact the supervisor Tim Bynum <linux-howto@sunsite.unc.edu>.

Tim Bynum will post the listing to several national and international newsgroups on a monthly basis. In addition, the Linux VAR HOWTO can be found on the World Wide Web at http://www.linuxports.com. New versions of the Linux VAR HOWTO are always placed at this site first, so please be sure to check if the copy you are reading is still up to date!

Companies selling hardware for the purpose of running Linux are invited to fill out the form located at . Companies offering other services should so a more appropriate template can be worked out.

2 Copyright Information

This HOWTO s Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999 Mr. Poet and LinuxPorts

A verbatim copy may be reproduced or distributed in any medium physical or electronic without permission of the author. Translations are similarly permitted without express permission if it includes a notice on who translated it.

Short quotes may be used without prior consent by the author. Derivative work and partial distributions of the Linux VAR HOWTO must be accompanied with either a verbatim copy of this file. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions. In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs. We further want that all information provided in the HOWTOs is disseminated. If you have questions, please contact Tim Bynum, the Linux HOWTO coordinator, at linux-howto@metalab.unc.edu.

3 VAR ID: Australia, Cybersource

     Level 8
     140 Queen Street
     Melbourne, NON United States 3000
     Australia

Platform: x86 AXP Distribution: RedHat Company Phone Number:

+61 3 9642 5997 Company Fax:

+61 3 9642 5997 Contact Email Address:

info@cyber.com.au Company URL:

http://www.cyber.com.au Primary Contact:

Con Zymaris

4 VAR ID: Australia, EMUSYS Unix Consulting

     20 Ivanhoe Rd
     Croydon
     Sydney, NON United States 2132
     Australia

Platform: x86 AXP PPC Distribution: Caldera Debian RedHat Suse LinuxPPC Company Phone Number:

+61 2 9798 7604 Company Fax:

+61 2 9798 2854 Contact Email Address:

sales@emusys.com.au Company URL:

http://www.emusys.com.au Primary Contact:

Anthony Rumble

5 VAR ID: Australia, Sydnet Group Pty Ltd

     PO Box 4664
     North Rocks, NON United States NSW 2151
     Australia

Platform: x86 AXP Sparc M68k Distribution: Caldera Debian RedHat LinuxRouter UltraPenguin Company Phone Number:

+61 2 9873 6400 Company Fax:

+61 2 9873 6411 Contact Email Address:

sales@syd.net.au Company URL:

http://www.syd.net.au Primary Contact:

Mark Lipscombe

6 VAR ID: Austria, QUANT-X Service & Consulting GmbH

     Hintnausdorf 14/1
     Sankt Veit a.d.Glan, NON United States
     Austria

Platform: AXP Distribution: RedHat Company Phone Number:

+43 4212 90555-0 Company Fax:

+43 4212 90555-20 Contact Email Address:

office@quant-x.com Company URL:

http://www.quant-x.com Primary Contact:

Wolfgang Dreyer

7 VAR ID: Brazil, Millennium Consultoria e Informática Ltda.

     R. Francisco Parolin, 758
     Curitiba, NON United States 802220-360
     Brazil

Platform: x86 Distribution: RedHat Company Phone Number:

+55-41-333-5645 Company Fax:

+55-41-333-5645 Contact Email Address:

millennium@millennium.etc.br Company URL:

http://www.millennium.etc.br Primary Contact:

André A. C. Santos

8 VAR ID: Canada, Affinity Systems Inc.

     Box 10, RR 1
     Innisfail, AB T4G 1T6
     Canada

Platform: x86 AXP Distribution: Caldera RedHat Company Phone Number:

1-403-886-4638 Company Fax:

1-403-886-2663 Contact Email Address:

sales@affinity-systems.ab.ca Company URL:

http://www.affinity-systems.ab.ca Primary Contact:

James Bourne

9 VAR ID: Canada, Centurion Services

     9351 Blundell Road
     Richmond, BC V6Y 1K5
     Canada

Platform: x86 AXP Distribution: RedHat Suse Company Phone Number:

604-279-1857 Company Fax:

604-279-1800 Contact Email Address:

jwalter@home.com Company URL:

http://www.ipipeline.net/centurion/ Primary Contact:

Mr.Jan Walter

10 VAR ID: Canada, Mediabase

     3 Shenston rd.
     look for new location soon!!
     North York, ON m2r 3b3
     Canada

Platform: x86 Distribution: Debian RedHat Company Phone Number:

416-468-5400 Company Fax:

no Contact Email Address:

mediabase@ihosts.net Company URL:

http://www.mediabase.ihosts.net Primary Contact:

Cyrus Hill

11 VAR ID: Canada, Net Direct Inc.

     557 Havelock Dr.
     Waterloo, ON N2L 4Z1
     Canada

Platform: x86 AXP Distribution: Caldera Debian RedHat Company Phone Number:

519-579-5006 Company Fax:

519-745-9940 Contact Email Address:

john@netdirect.ca Company URL:

http://www.netdirect.ca Primary Contact:

John Van Ostrand

12 VAR ID: Canada, Softouch Computer Solutions

     5151 3rd Street SE
     Calgary, AB T2H2X6
     Canada

Platform: x86 Distribution: RedHat Company Phone Number:

403 Company Fax:

258-0267 Contact Email Address:

sales@softouchcomputer.com Company URL:

http://www.softouchcomputer.com Primary Contact:

Al Green

13 VAR ID: Canada, VL

     221 Britannia
     Ottawa, ON K2B 5X1
     Canada

Platform: x86 Distribution: RedHat Company Phone Number:

1 (613) 292-9818 Company Fax:

1 (613) 829-0967 Contact Email Address:

hungvu@netcom.ca Company URL:

http:// Primary Contact:

Hung Vu

14 VAR ID: Finland, SOT Finnish Software Engineering Ltd.

     Hermiankatu 8 E
     FIN-33720
     TAMPERE, NON United States
     Finland

Platform: x86 Distribution: RedHat Company Phone Number:

+358-3-316 5544 Company Fax:

+358-3-316 5959 Contact Email Address:

linux@sot.com Company URL:

http://www.bestlinux.net Primary Contact:

Kai Valijärvi

15 VAR ID: France, Paralline

     71, avenue des Vosges
     -
     STRASBOURG, NON United States 67000
     France

Platform: x86 Distribution: Suse Company Phone Number:

33 3 88 14 17 40 Company Fax:

33 3 88 14 17 41 Contact Email Address:

sales@paralline.com Company URL:

http://www.paralline.com Primary Contact:

Pierre BRUA

16 VAR ID: Germany, IGEL GmbH

     Aindlinger Str. 16
     Augsburg, NON United States 86167
     Germany

Platform: x86 Distribution: Suse Company Phone Number:

+49-821-7200-200 Company Fax:

+49-821-7200-189 Contact Email Address:

sales@igel.de Company URL:

http://www.igel.de Primary Contact:

Mike Finckh

17 VAR ID: Germany, IGEL

     Aindlinger Str. 16
     Augsburg, NON United States 86167
     Germany

Platform: x86 Distribution: Suse Company Phone Number:

+49-821-7200-200 Company Fax:

+49-821-7200-189 Contact Email Address:

sales@igel.de Company URL:

http://www.igel.de Primary Contact:

Mike Finckh

18 VAR ID: Germany, dynamis EDV-Consulting

     Bundschuhstrasse 24
     76661 Philippsburg, NON United States
     Germany

Platform: x86 AXP Distribution: Caldera RedHat Suse Company Phone Number:

+49 7256 9326-0 Company Fax:

+49 7256 9326-90 Contact Email Address:

info@dynamis.de Company URL:

http://www.dynamis.de Primary Contact:

Jochen Lillich

19 VAR ID: Hong Kong, Genesis Systems Int’l Ltd.

     5/F., 7/F., Honytex Building
     22 Ashley Road, T.S.T.
     Kowloon, NON United States
     Hong Kong

Platform: AXP Distribution: RedHat Company Phone Number:

+852 2815 0728 Company Fax:

+852 2815 0729 Contact Email Address:

info@genesis.com.hk Company URL:

http://www.genesis.com.hk Primary Contact:

Gyver Lo

20 VAR ID: India, G.T.Enterprises

     # 912, 14th Main, 4th Cross
     Maruthi Circle, Hanumanthanagar
     Bangalore, NON United States 560019
     India

Platform: x86 Distribution: RedHat Company Phone Number:

+91-80-6606093 Company Fax:

+91-80-6671407 Contact Email Address:

gtcdrom@vsnl.com Company URL:

http://www.gtcdrom.com Primary Contact:

U.M. Taranath

21 VAR ID: Indonesia, PT Cakram DataLingga Duaribu (CDL2000)

     Gunung Batu 131
     Bogor, NON United States 16610
     Indonesia

Platform: x86 Distribution: RedHat Company Phone Number:

+62-251-332122 Company Fax:

+62-251-332122 Contact Email Address:

info@cdl2000.or.id Company URL:

http://www.cdl2000.or.id Primary Contact:

Maliana Harsoadi

22 VAR ID: Italy, Conecta Telematica srl

     Viale Tricesimo 101/13
     Udine, NON United States 33100
     Italy

Platform: x86 AXP PPC Distribution: Debian RedHat Suse LinuxPPC Company Phone Number:

+39 432 548825 Company Fax:

+39 432 299141 Contact Email Address:

info@conecta.it Company URL:

http://www.conecta.it/linux Primary Contact:

Carlo Daffara

23 VAR ID: Italy, Nauta Srl - Internet per l’impresa

     Viale Tricesimo 181
     Udine (UD), NON United States I-33100
     Italy

Platform: x86 Distribution: Debian RedHat LinuxRouter Company Phone Number:

+39 0432 477588 Company Fax:

+39 0432 548314 Contact Email Address:

info@Nauta.it Company URL:

http://www.Nauta.it Primary Contact:

Antonio Pezzoni

24 VAR ID: Italy, Studio C

     via Romeo Romei 23
     Rome, NON United States 00136
     Italy

Platform: x86 AXP Sparc PPC M68k Distribution: Caldera Debian RedHat Suse LinuxPPC LinuxRouter UltraPenguin Company Phone Number:

+39639722036 Company Fax:

+39639736598 Contact Email Address:

info@studioc-advertising.com Company URL:

http://www.studioc-advertising.com Primary Contact:

Marco Capecci

25 VAR ID: Japan, SoftWare Research Co.

     Hayabusacho 2-11
     Chiyodaku
     Tokyo, NON United States 102-0092
     Japan

Platform: x86 Distribution: Caldera Debian RedHat Company Phone Number:

81-3-3264-6171 Company Fax:

81-3-3264-0647 Contact Email Address:

info@src.co.jp Company URL:

http://www.src.co.jp Primary Contact:

Yossihiro Hayama

26 VAR ID: Malaysia, Linux Resources Sdn. Bhd.

     No. 33-1, Jalan 46A/26
     Pusat Bandar Taman Sri Rampai
     Setapak, NON United States 54000
     Malaysia

Platform: x86 AXP Sparc PPC ARM M68k Distribution: Caldera Debian RedHat Suse LinuxPPC LinuxRouter UltraPenguin Company Phone Number:

+603-4139010 Company Fax:

+603-4139011 Contact Email Address:

root@ihsan.com Company URL:

http://linux.ihsan.com Primary Contact:

Liwauddin Mohamad

27 VAR ID: Malaysia, MAGNIFIX

     2102 Wisma Rampai
     Jln 34/26, Taman Seri Rampai, Setapak
     Kuala Lumpur, NON United States 53300
     Malaysia

Platform: x86 Distribution: RedHat Company Phone Number:

603 412 1775 Company Fax:

603 412 1550 Contact Email Address:

izauddin@magnifix.com.my Company URL:

http://www.magnifix.com.my Primary Contact:

Izauddin Isa

28 VAR ID: Malaysia, Malaysian Linux Competency Center Sdn. Bhd.

     32-A, Jalan Mamanda 7,
     Ampang Point
     Ampang, NON United States 68000 Selangor
     Malaysia

Platform: x86 Sparc Distribution: RedHat UltraPenguin Company Phone Number:

+60 3 4571798 Company Fax:

+60 3 4523669 Contact Email Address:

linux@mlcc.com.my Company URL:

http://www.mlcc.com.my Primary Contact:

Abd Rahman Johari

29 VAR ID: Netherlands, CMG Trade, Transport & Industry

     Kralingseweg 241-249
     Postbus 8566
     Rotterdam, NON United States 3009 AN
     Netherlands

Platform: x86 AXP Sparc PPC ARM M68k Distribution: Caldera Debian RedHat Suse LinuxPPC LinuxRouter UltraPenguin Company Phone Number:

+31.10.2537000 Company Fax:

+31.10.2537032 Contact Email Address:

Henk.Drenth@cmg.nl Company URL:

http://www.cmg.nl Primary Contact:

Henk Drenth

30 VAR ID: Netherlands, X OS Experts in Open Systems BV

     Kruislaan 419
     Amsterdam, NON United States 1098 VA
     Netherlands

Platform: x86 AXP Sparc Distribution: Debian RedHat Suse Company Phone Number:

+31 20 6938364 Company Fax:

+31 20 6948204 Contact Email Address:

info@xos.nl Company URL:

http://www.xos.nl Primary Contact:

Jos Vos

31 VAR ID: Poland, Emit Sp z o.o.

     ul. Koszarowa 48/9
     Wroclaw, NON United States
     Poland

Platform: x86 Distribution: Company Phone Number:

+48713267026 Company Fax:

+48723289733 Contact Email Address:

kontakt@emit.pl Company URL:

http://www.emit.pl Primary Contact:

Ian Carr-de Avelon

32 VAR ID: Romania, EVERCOMM Serv SRL

     Str. Ciucea 8, bl.L18, ap.189
     sector 3
     Bucharest, NON United States 7400
     Romania

Platform: x86 Distribution: RedHat Company Phone Number:

+40-9-254-4355 Company Fax:

+40-1-345-4355 Contact Email Address:

Mihnea.Stoenescu@evercomm.ro Company URL:

http://www.evercomm.ro Primary Contact:

Mihnea Stoenescu

33 VAR ID: Spain, Innovacion y Futuro s.l.

     Plaza de S'Antigor, 3
     Manacor, NON United States 07500
     Spain

Platform: x86 Distribution: Caldera Company Phone Number:

+34 971846014 Company Fax:

+34 971846014 Contact Email Address:

info@i-futuro.com Company URL:

http://www.i-futuro.com Primary Contact:

Pascal Peregrina

34 VAR ID: Spain, LANGRE INGENIEROS S.L.

     Castilla 95.
     Santander., NON United States 39009
     Spain

Platform: x86 Distribution: Suse Company Phone Number:

34-942-33-56-33 Company Fax:

34-942-35-42-75 Contact Email Address:

arturo@langre.com Company URL:

http://www.langre.com Primary Contact:

Arturo Limon

35 VAR ID: Sweden, South Pole AB

     Banvaktsvägen 12
     Solna, NON United States 171 48
     Sweden

Platform: x86 AXP Distribution: Debian RedHat Suse Company Phone Number:

+46 (0)8 566 106 50 Company Fax:

+46 (0)8 566 106 50 Contact Email Address:

info@southpole.se Company URL:

http://southpole.se Primary Contact:

Jakob Sandgren

36 VAR ID: Taiwan, Whinstone Info. Corp.

     4f, No 15, Lane 2
     Dong Shing Rd.
     Taipei, NON United States 105
     Taiwan

Platform: x86 AXP Sparc Distribution: RedHat Company Phone Number:

886-2-27492626 Company Fax:

886-2-27468790 Contact Email Address:

info@whinstone.com Company URL:

http://www.whinstone.com Primary Contact:

Anthony Hsia

37 VAR ID: Turkey, Gelecek Bilisim ve Iletisim A.S

     Perpa Ticaret Merkezi B Blok
     Kat 11 No:1932
     Istanbul, NON United States
     Turkey

Platform: x86 Distribution: Caldera Debian RedHat Suse Company Phone Number:

+90 212 2107194 Company Fax:

+90 312 2100442 Contact Email Address:

info@gelecek.com.tr Company URL:

http://www.gelecek.com.tr Primary Contact:

Gorkem Cetin

38 VAR ID: US, AZ, Sunset Systems

     1328 East Barbarita Avenue
     Gilbert, AZ 85234
     US

Platform: x86 Distribution: RedHat Company Phone Number:

602-770-3161 Company Fax:

480-497-8012 Contact Email Address:

sales@sunsetsystems.com Company URL:

http://www.sunsetsystems.com Primary Contact:

Rod Roark

39 VAR ID: US, CA, ASL Workstations, Inc.

     1757 Houret Court
     Milpitas, CA 95035
     US

Platform: x86 Distribution: Caldera RedHat Suse Company Phone Number:

877 ASL 3535 Company Fax:

408 941 2071 Contact Email Address:

sales@aslab.com Company URL:

http://www.aslab.com Primary Contact:

Jim Truong

40 VAR ID: US, CA, Linux Laptops Ltd.

     3583 Lynx Dr.
     San Jose, CA 95136
     US

Platform: x86 Distribution: Debian Company Phone Number:

+1 888 546-8958 Company Fax:

+1 408 269 9872 Contact Email Address:

info@linuxlaptops.com Company URL:

http://www.linuxlaptops.com Primary Contact:

Nathan Myers

41 VAR ID: US, CA, PSSC Labs

     27601 Forbes Rd Ste# 59
     Laguna Niguel, CA 92677
     US

Platform: x86 AXP Distribution: Caldera RedHat Company Phone Number:

949-582-7088 Company Fax:

949-582-7093 Contact Email Address:

4sales@pssclabs.com Company URL:

http://www.pssclabs.com Primary Contact:

Jason Smith

42 VAR ID: US, CO, Apache Digital Corporation

     777 Main Avenue
     #211
     Durango, CO 81301
     US

Platform: x86 AXP Distribution: Caldera Debian RedHat Suse Company Phone Number:

970-259-8153 Company Fax:

970-259-8145 Contact Email Address:

info@apache.com Company URL:

http://www.apache.com Primary Contact:

Steve Hauquitz

43 VAR ID: US, CT, GNUTEC

     1655 North Colony Road, Suite 3002
     Meriden, CT 06450
     US

Platform: x86 AXP Sparc PPC ARM M68k Distribution: Caldera Debian RedHat Suse LinuxPPC LinuxRouter UltraPenguin Company Phone Number:

1-203-379-0125 Company Fax:

1-203-379-0126 Contact Email Address:

info@gnutec.com Company URL:

http://www.gnutec.com Primary Contact:

Kyle Amon

44 VAR ID: US, FL, eBuilders Network

     8204 Crystal Clear Lane
     Suite 1200
     Orlando, FL 32809
     US

Platform: x86 Distribution: Caldera RedHat Company Phone Number:

407-857-6800 Company Fax:

407-857-2333 Contact Email Address:

pablo@ebuildernet.com Company URL:

http://www.ebuildernet.com Primary Contact:

Pablo Bryan

45 VAR ID: US, IL, Advanced Horizons, Inc

     323 Ferndale Ave
     Elmhurst, IL
     US

Platform: x86 Distribution: Suse Company Phone Number:

630-941-9332 Company Fax:

630-941-9338 Contact Email Address:

roger@ahinc.com Company URL:

http://www.ahinc.com Primary Contact:

Roger Petersen

46 VAR ID: US, IL, QLITech Linux COmputers

     3717 43rd st SUite #112
     Moline, IL 61265
     US

Platform: x86 Distribution: Caldera Debian RedHat Company Phone Number:

1-877-24-linux Company Fax: Contact Email Address:

sales@qlitech.net Company URL:

http://www.qlitech.net Primary Contact:

Ray Sanders

47 VAR ID: US, IL, QLITech Linux Computers

     3717 43rd st. Suite #112
     Moline, IL 61265
     US

Platform: x86 Distribution: Caldera Debian RedHat Suse Company Phone Number:

1-877-24-LINUX Company Fax:

1-309-764-1142 Contact Email Address:

sales@qlitech.net Company URL:

http://www.qlitech.net Primary Contact:

Ray Sanders

48 VAR ID: US, IL, WellThot Inc.

     305 Van Damin Ave.
     Glen Ellyn, IL 60137-5215
     US

Platform: x86 Distribution: RedHat LinuxRouter Company Phone Number:

630-545-1470 Company Fax:

630-545-1471 Contact Email Address:

sales@wellthot.com Company URL:

http://wellthot.com Primary Contact:

Mark Schwenk

49 VAR ID: US, KY, IndyBox Systems, Inc.

     PO Box 3564
     Carmel, KY 46082-3564
     US

Platform: x86 AXP Sparc PPC ARM M68k Distribution: Caldera Debian RedHat Suse LinuxPPC LinuxRouter UltraPenguin Company Phone Number:

+01 (317) 846.9762 Company Fax: Contact Email Address:

info@indybox.com Company URL:

http://www.indybox.com Primary Contact:

Charles Martin

50 VAR ID: US, KY, Neville Technologies

     4240 Frankfort Road
     Shelbyville, KY 40065
     US

Platform: x86 Distribution: RedHat Company Phone Number:

502-633-3711 Company Fax:

502-633-7950 Contact Email Address:

info@neville.net Company URL:

http://www.neville.net Primary Contact:

Chuck Neville

51 VAR ID: US, MA, Network Concierge Inc.

     350 Pleasant St.
     Ashland, MA 01721
     US

Platform: x86 Distribution: Company Phone Number:

1- 877- 876-1169 Company Fax:

1-508-881-7374 Contact Email Address:

sales@nc4u.com Company URL:

http://www.nc4u.com Primary Contact:

Leonardo Salazar Sr.

52 VAR ID: US, MD, AmNet Computers

     3618 Anne Hasthaway
     Suite 1B
     Randallstown, MD 21133
     US

Platform: x86 AXP Distribution: RedHat Company Phone Number:

1 410 922-3554 Company Fax:

1 410 922-3552 Contact Email Address:

amnet@amnet-comp.com Company URL:

http://amnet-comp.com Primary Contact:

Joseph Volodarsky

53 VAR ID: US, MN, Cranston Software

     3255 Hennepin Av.S.
     Minneapolis, MN 55408
     US

Platform: x86 Distribution: Debian Company Phone Number:

612 Company Fax:

378-9052 Contact Email Address:

sales@cranston.com Company URL:

http://cranston.com Primary Contact:

M.A. al-Kabir

54 VAR ID: US, MN, Donnal Group Inc

     19820 Cottagewood Road
     Excelsior, MN 55331
     US

Platform: x86 Distribution: Caldera Company Phone Number:

612-474-7785 Company Fax:

612-474-7793 Contact Email Address:

info@donnalgroup.com Company URL:

http://www.donnalgroup.com Primary Contact:

Tim Burke

55 VAR ID: US, MO, Atipa Linux Solutions

     2608 S. Main
     Joplin, MO 64804
     US

Platform: x86 Distribution: Caldera RedHat Company Phone Number:

(800) 360-4346 Company Fax:

(417) 626-2692 Contact Email Address:

info@atipa.com Company URL:

http://www.atipa.com Primary Contact:

Brandon Fuhr

56 VAR ID: US, MO, Omnitec Corporatin

     PO Box 411010
     St. Louis, MO 63141
     US

Platform: x86 Distribution: Debian RedHat Suse LinuxRouter Company Phone Number:

3149974600 Company Fax:

3149970153 Contact Email Address:

info@omnitec.net Company URL:

http://www.omnitec.net Primary Contact:

Lee Lammert

57 VAR ID: US, NC, ApeX Systems Integration Corp.

     PO Box 338
     Apex, NC 27502
     US

Platform: x86 Distribution: RedHat Suse Company Phone Number:

919-468-8150 Company Fax:

919-468-5288 Contact Email Address:

apex@2boot.com Company URL:

http://www.2boot.com Primary Contact:

Joe Ferguson

58 VAR ID: US, NC, ComputerNet Inc.

     8320 Univ Exec PK Dr
     Suite 105
     Charlotte, NC 28262
     US

Platform: x86 Distribution: RedHat Company Phone Number:

704-510-5292 Company Fax:

704-510-5296 Contact Email Address:

kbrust@computernetinc.com Company URL:

http://http://www.computernetinc.com Primary Contact:

Kurt Brust

59 VAR ID: US, NC, Indelible Blue

     3209 Gresham Lake Road
     Suite 135
     Raleigh, NC 27615
     US

Platform: x86 Distribution: Caldera Debian RedHat Suse LinuxRouter Company Phone Number:

1.919.878.9700 or 1.800.776.8284 Company Fax:

1.919.878.7479 Contact Email Address:

sales@indelible-blue.com Company URL:

http://www.indelible-blue.com Primary Contact:

Mark Turner

60 VAR ID: US, NE, BrainStorm Technologies

     1525 GrandView Ste#1
     Papillion, NE
     US

Platform: x86 AXP PPC Distribution: RedHat Company Phone Number:

1-800-598-6732 Company Fax:

1-800-598-6732 Contact Email Address:

info@bstc.net Company URL:

http://www.bstc.net Primary Contact:

Brian Roberson

61 VAR ID: US, NE, BrainStorm Technologies

     1525 GrandView Ste#1
     Papillion, NE
     US

Platform: x86 AXP PPC Distribution: RedHat Company Phone Number:

1-800-598-6732 Company Fax:

1-800-598-6732 Contact Email Address:

info@bstc.net Company URL:

http://www.bstc.net Primary Contact:

Brian Roberson

62 VAR ID: US, NH, DCG Computers, Inc.

     4 Sanborn Rd., # 2
     Londonderry, NH 03053
     US

Platform: x86 AXP Sparc Distribution: Caldera Debian RedHat Suse Company Phone Number:

+1-603-421-1800 Company Fax:

+1-603-421-0911 Contact Email Address:

sjg@dcginc.com Company URL:

http://www.dcginc.com Primary Contact:

Stephen Gaudet

63 VAR ID: US, NON United States, Microway

     Research Park
     Box 79
     Kingston Ma, NON United States 02364
     US

Platform: AXP Distribution: RedHat Company Phone Number:

508-746-7341 Company Fax:

508-746-4678 Contact Email Address:

info@microway.com Company URL:

http://www.microway.com Primary Contact:

Ann Fried

64 VAR ID: US, NON United States, Telenet System Solutions, Inc.

     2480 Kruse Drive
     San Jose, NON United States
     US

Platform: x86 Distribution: RedHat Company Phone Number:

1-877-BootLinux Company Fax:

408-383-0334 Contact Email Address:

linux@tesys.com Company URL:

http://www.tesys.com Primary Contact:

Kamal Singh

65 VAR ID: US, NY, Logicept Corporation

     41 Chenango Street, Suite 201
     Binghamton, NY 13901-2901
     US

Platform: x86 Distribution: RedHat Company Phone Number:

888.875.1355 Company Fax: Contact Email Address:

info@logicept.com Company URL:

http://www.logicept.com Primary Contact:

Jaron Rubenstein

66 VAR ID: US, OH, 800Linux

     1357 W. Lane Ave suite 211
     Columbus, OH 43221
     US

Platform: x86 Distribution: RedHat Company Phone Number:

614-324-1610 Company Fax:

614-324-1617 Contact Email Address:

sales@gtltd.com Company URL:

http://800linux.com Primary Contact:

67 VAR ID: US, OH, LANshark Systems, Inc.

     784 Morrison Road
     Columbus, OH 43230
     US

Platform: x86 Distribution: Caldera Debian RedHat Suse Company Phone Number:

614-751-1111 Company Fax:

614-751-1112 Contact Email Address:

info@lanshark.com Company URL:

http://linux.lanshark.com Primary Contact:

Scott Sharkey

68 VAR ID: US, OH, The Computer Underground

     1357 West Lane Ave suite 210
     Columbus, OH
     US

Platform: x86 Distribution: Caldera Debian RedHat Suse Company Phone Number:

614-485-0506 Company Fax: Contact Email Address:

sales@tcu-inc.com Company URL:

http://www.tcu-inc.com Primary Contact:

Mark Nielsen

69 VAR ID: US, OH, Unique Systems, Inc.

     6920 Spring Valley Dr.
     Suite 106
     Holland, OH 43528
     US

Platform: x86 AXP Distribution: Caldera RedHat Suse Company Phone Number:

419 861-3331 Company Fax:

419 861-3340 Contact Email Address:

info@uniqsys.com Company URL:

http://www.uniqsys.com Primary Contact:

Glenn Jacobson

70 VAR ID: US, OR, 3D Technology, Inc.

     9150 SW Pioneer Ct.,
     Suite E
     Wilsonville , OR 97070
     US

Platform: x86 Distribution: Caldera RedHat Company Phone Number:

1-503-582-1190 Company Fax:

1-503-582-9102 Contact Email Address:

netsales@3dtech.net Company URL:

http://www.3dtech.net Primary Contact:

Kerry Lafferty

71 VAR ID: US, PR, OG Consulting

     Urb. Santa Clara
     W-1 Anamu
     Guaynabo, PR 00969
     US

Platform: x86 Distribution: Caldera RedHat Company Phone Number:

787-731-4332 Company Fax:

787-731-4331 Contact Email Address:

james@og-consulting.com Company URL:

http://www.og-consulting.com Primary Contact:

James O’Malley

72 VAR ID: US, TX, Computer Innovations International

     2145 NW Military Highway, Suite 103
     San Antonio, TX 78213
     US

Platform: x86 AXP Sparc Distribution: Caldera RedHat Suse LinuxRouter Company Phone Number:

210-348-0850 Company Fax:

210-348-0851 Contact Email Address:

wvega@txcii.com Company URL:

http://www.txcii.com Primary Contact:

Waldemar Vega

73 VAR ID: US, TX, Heights Open Systems

     517 Harvard
     Houston, TX 77007
     US

Platform: x86 AXP Distribution: Debian RedHat LinuxRouter Company Phone Number:

713-864-5676 Company Fax:

713-864-5676 Contact Email Address:

perry@open-systems.com Company URL:

http://www.open-systems.com Primary Contact:

Perry Piplani

74 VAR ID: US, TX, Parachute

     3500 Oakmont
     Suite 101
     Austin, TX 78731
     US

Platform: x86 Distribution: RedHat LinuxRouter Company Phone Number:

(512)374-1898 Company Fax:

(512)374-9753 Contact Email Address:

kjones@parachute.com Company URL:

http://www.parachute.com Primary Contact:

Kristian Jones

75 VAR ID: US, VA, NetSource of Virginia, Inc.

     7639 Hull Street Rd.
     Suite 202
     Richmond, VA 23235-6438
     US

Platform: x86 Distribution: RedHat Company Phone Number:

1-804-674-6040 Company Fax:

1-804-674-6041 Contact Email Address:

netsource@erols.com Company URL:

http://www.erols.com/netsource Primary Contact:

Jeff Davis

76 VAR ID: US, WA, Cheek Consulting

     620 5th Ave W
     Suite 309
     Seattle, WA 98119
     US

Platform: x86 Distribution: Caldera Company Phone Number:

+1 (206) 282-2892 Company Fax:

+1 (206) 727-6797 Contact Email Address:

joseph@cheek.com Company URL:

http://www.cheek.com Primary Contact:

Joseph Cheek

77 VAR ID: US, WA, Lodgepole Technology

     14578 NE 177th Drive
     n/a
     Redmond, WA 98072
     US

Platform: AXP Distribution: RedHat Company Phone Number:

1(425)867-9091 Company Fax:

1(425)558-5752 Contact Email Address:

info@lodgepole.com Company URL:

http://www.lodgepole.com Primary Contact:

Geordy Rostad

78 VAR ID: United Arab Emirates, JM Infotech Solutions STPI

     P.O. Box 692
     AL MOTAMEM
     Karama, NON United States 692
     United Arab Emirates

Platform: x86 Distribution: RedHat Company Phone Number:

971 50 6326822 Company Fax:

971 4 359295 Contact Email Address:

mammen@emirates.net.ae Company URL:

http://linuxsolutions.hypermart.net Primary Contact:

John Mammen

79 VAR ID: United Kingdom, DGC-NMS

     7 Owen Road
     Rainhill, Prescot
     Merseyside, NON United States L35 0PJ
     United Kingdom

Platform: x86 AXP Sparc PPC M68k Distribution: Caldera Debian RedHat Suse LinuxPPC Company Phone Number:

+44 151 289 9250 Company Fax:

+44 151 289 7895 Contact Email Address:

rich@dgc-nms.co.uk Company URL:

http://www.dgc-nms.co.uk Primary Contact:

Richard Bradshaw

80 VAR ID: United Kingdom, Definite Linux Systems Ltd

     264 Ripon Road
     Stevenage, NON United States SG1 4NF
     United Kingdom

Platform: x86 Distribution: Caldera Debian RedHat Suse LinuxRouter Company Phone Number:

+44 1438 223433 Company Fax: Contact Email Address:

sales@definitelinux.com Company URL:

http://www.definitelinux.com Primary Contact:

Jason Clifford

81 VAR ID: United Kingdom, Digital Networks

     Unit 12, New Mill
     Park Road, Dukinfield
     Cheshire, NON United States SK16 5LX
     United Kingdom

Platform: x86 Distribution: Caldera RedHat Suse Company Phone Number:

44 161 339 8555 Company Fax:

44 161 339 8666 Contact Email Address:

info@dnuk.com Company URL:

http://dnuk.com Primary Contact:

Lee Chisnall

82 VAR ID: United Kingdom, iCode Systems Ltd

     Consort Court
     High Street
     Fareham, NON United States PO16 7AL
     United Kingdom

Platform: x86 Distribution: Caldera Debian RedHat Suse Company Phone Number:

+44 1329 829898 Company Fax:

+44 1329 829899 Contact Email Address:

linux@icode.co.uk Company URL:

http://www.icode.co.uk Primary Contact:

Darren Durbin